Grade 11Trigonometry

Given : Sec theta+ tan theta =3 , where theta belongs to (0,π/2). Find cos theta

Isha Keshav
8 Years agoGrade 11

1 Answer

Arun
ApprovedApproved Tutor Answer8 Years ago
Dear Isha
 
As we know
sec²@ - tan²@ = 1
(Sec@ + tan@)(sec@ - tan@) = 1
Hence sec@ - tan@ = 1/3
Now
2sec @ = 3 + 1/3 = 10/3
sec @ = 5/3
Hence cos @ = 3/5
In case of any difficulty, please feel free to ask.
Regards
Arun (askIITIans forum expert)